3. The function f is defined by f(x) = ax + b for x € R, where a and b are constants. It



is given that f(2) = 1 and f(5) = 7.



i) Set up a pair of simultaneous equations using the information given.



ii) Solve them and find the values of a and b.



iii) Solve the equation f(f(x)) = 0.


1
Expert's answer
2021-11-18T17:24:26-0500

i)

"2a+b=1"

"5a+b=7"


ii)

"3a=6"

"a=2,b=1-2a=1-4=-3"

"f(x)=2x-3"


iii)

"f(f(x)) =2(2x-3)-3"

"2(2x-3)-3=4x-9=0"

"x=9\/4=2.25"